# ID-Booth: Identity-consistent Face Generation with Diffusion Models

### v3

@@ V Conclusion
In this paper, we presented ID-Booth, a new diffusion-based fine-tuning framework for generating high-quality identity-consistent images, suitable for augmenting existing small-scale datasets in a privacy-preserving manner. To this end, ID-Booth relies on a novel triplet identity training objective that improves both intra-identity consistency and inter-identity separability, while retaining image diversity of pretrained state-of-the-art diffusion models. Throughout the experiments, we showcase that training deep recognition models on datasets augmented with synthetic samples of ID-Booth results in better performance across five verification benchmarks than when performing augmentation with synthetic samples of existing approaches or training only on real-world data. With regards to future work, we aim to investigate the applicability of identity-based objectives in the training of conditioning approaches and exploring the creation of larger-scale datasets.
VI Acknowledgements
Supported in parts by the Slovenian Research and Innovation Agency (ARIS) through the Research Programmes P2-0250 (B) ”Metrology and Biometric Systems” and P2–0214 (A) “Computer Vision”, the ARIS Project J2-50065 ”DeepFake DAD”, and the ARIS Young Researcher Programme.
@@ VI Acknowledgements
VI-A Additional results and ablation studies
@@
VI-B Details regarding the evaluation methodology

### v4

@@ V Conclusion
In this paper, we presented ID-Booth, a new diffusion-based fine-tuning framework for generating high-quality identity-consistent face images, suitable for augmenting existing small-scale recognition datasets in a privacy-preserving manner. To this end, ID-Booth relies on a novel triplet identity training objective that improves both intra-identity consistency and inter-identity separability, while retaining image diversity of pretrained state-of-the-art diffusion models. Throughout the experiments, we showcase that training deep recognition models on datasets augmented with synthetic samples of ID-Booth results in better performance across five verification benchmarks than when performing augmentation with synthetic samples of existing approaches or training only on real-world data. With regard to future work, we aim to investigate the applicability of identity-based objectives in the training of conditioning approaches and plan to explore the creation of larger-scale datasets.
@@
V-A Additional results and ablation studies
@@
V-B Details regarding the evaluation methodology